# Estimation of uncertainties in the density driven flow in fractured porous media using MLMC

### v1

@@ Abstract
We use the Multi Level Monte Carlo method to estimate uncertainties in a Henry-like salt water intrusion problem with a fracture. The flow is induced by the variation of the density of the fluid phase, which depends on the mass fraction of salt. We assume that the fracture has a known fixed location but an uncertain aperture. Other input uncertainties are the porosity and permeability fields and the recharge. In our setting, porosity and permeability vary spatially and recharge is time-dependent. For each realisation of these uncertain parameters, the evolution of the mass fraction and pressure fields is modelled by a system of non-linear and time-dependent PDEs with a jump of the solution at the fracture. The uncertainties propagate into the distribution of the salt concentration, which is an important characteristic of the quality of water resources. We show that the multilevel Monte Carlo (MLMC) method is able to reduce the overall computational cost compared to classical Monte Carlo methods. This is achieved by balancing discretisation and statistical errors. Multiple scenarios are evaluated at different spatial and temporal mesh levels. The deterministic solver ug4 is run in parallel to calculate all stochastic scenarios.
@@ 1 Introduction
An important and challenging problem in hydrogeology is the modeling of densiity-driven subsurface flow and salt transport in fractured aquifers. Essential difficulties in these settings are complicated and parameter-sensitive flow pattern due to the fractures as well as a number of the uncertain parameters. Uncertainties in the fracture geometry, porosity, permeability, recharge can strongly influence the evolution of the salt concentration in the subsurface (cf. [1, 3, 4, 33, 44]). In a real reservoir the number of the fractures, their location and their aperture are uncertain. One possible solution approach would be to model all these uncertainties, build a PCE-based surrogate, perform a sensitive analysis (see [5, 45]) to understand which uncertainties are very critical and which are not, and estimate how they contribute to the solution. This is a very non-trivial approach. Instead, we apply the Multi Level Monte Carlo method (MLMC) method. Another reason for the usage of MLMC method is that the deterministic problem is resolved via the multigrid method. So, there is a natural idea to try to couple the MLMC and multigrid methods.
An essential problem in water production from coastal aquifers is their salinization due to the intrusion of the seawater. Salt concentration in the drinking water should be kept below the prescribed limits, and therefore its forecast ois important for water resources management in coastal regions.
Fractures introduce heterogeneity in the aquifer that affects fluid flow and makes it difficult to predict its main characteristics. A challenge is that fractures can get sealed over time, reducing their permeability and affecting fluid flow. In summary, adding fractures to subsurface aquifers presents challenges related to complexity, sealing, heterogeneity and environmental concerns. However, with careful planning and management, fractures can also provide opportunities for increased production and improved reservoir performance.
As a model problem, we consider the benchmark from [20] which is a generalization of the Henry problem first introduced in [25]. The Henry problem became a benchmark for numerical solvers for the density-driven groundwater flow (see [50, 43, 42, 14]. In [39], the authors use the generalised polynomial chaos expansion approximation to investigate how incomplete knowledge of system properties affects the assessment of global quantities. In particular, they estimated the propagation of input uncertainties into a few dimensionless scalar parameters.
@@
In [20, 21, 38], the authors modeled and computed the density-driven flow in fractured porous media with deterministic settings. The fractures considered were represented by a $(d-1)$-dimensional manifolds in a $d$-dimensional domain. They developed a special numerical technique based on the finite volume method and implemented it in the ug3 library. In our present work, we apply the same model, the same discretization and solvers to compute the multiple uncertain scenarios.
In [21], the validity of the low-dimensional representation of the fractures has been addressed. The authors considered both $d$ and $(d-1)$-dimensional fractures. They have shown numerically that the $(d-1)$-dimensional representation can be a valid alternative to the full-dimensional resolution of all the subdomains provided that the fracture apertures are in some reasonable range. All the settings in this work were deterministic.
Other methods based on $(d-1)$-dimensional representation of fractures were considered in [1, 3, 4, 33, 44, 41, 19].
@@
Many techniques can be used to quantify uncertainties. One classical method is Monte Carlo (MC) sampling. It has a well-known disadvantage — a slow convergence $\mathcal{O}(\frac{1}{\sqrt{N}})$. Other relative recent techniques such as surrogate models and stochastic collocation may require a few hundred time-consuming simulations and assume a certain smoothness of the quantity of interest (QoI).
Another class of methods is the class of perturbation methods [12]. The idea is to decompose the QoI with respect to random parameters in a Taylor series. The higher order terms can be neglected for small perturbations, simplifying the analysis and numerics. These methods assume that the random perturbations are small. For larger perturbations, these methods usually do not work.
There are a number of studies where authors have modeled uncertainties in reservoirs (see [7, 48]). The link between stochastic methods and hydrogeological applications was made in [6]. Further the authors made recommendations to water suppliers in Germany about optimisation and risk assessment. The basics of stochastic hydrogeology and an overview of stochastic tools and uncertainty management are described in [40].
The review [46] deals with hydrogeological applications of recent advances in uncertainty quantification, probabilistic risk assessment and decision making under uncertainty.
@@
This work has the following structure. Section 2 describes the model. Section 3 contains description of numerical methods. In particular, the well-known MLMC method is reviewed in Section 3.2. Section 4 details the numerical results, which include the numerical analysis of the model problem, the computation of different statistics, the performance of the MLMC method, and the performance of the parallel solver with uncertain coefficients. Finally, we conclude this work with a discussion in Section 5.
@@ 2 Modeling and Problem Settings
We consider the density driven flow of the liquid phase (salt solution in water) with variable mass fraction of the salt in a fractured aquifer. The fractures are geological formations whose size in one of the geometrical directions is negligible in comparison with the scales of the domain whereas the permeability of the filling material is much higher than that of the surrounding medium. This causes essential problems when these fractures are considered as full-dimensional subdomains in the simulations: Specific, very fine grids needed for the resolution of this sort of the objects as well as the strong jump of the model parameters reduces the efficiency of the numerical methods significantly. For this reason, several other approaches have been developed where the fractures are represented by low-dimensional manifolds and the fracture aperture (width) is a parameter (cf. [1, 4, 16, 13, 32, 28, 38]). Whereas it has been shown that for the density-driven flow, these models provide a proper approximation only for a certain range of the fracture width ([21]), their application is considered to be reasonable in many situations. A particular advantage of such the models for the sampling methods is the possibility to change the fracture width without remeshing of the entire domain. In this paper, we assume that the porous matrix is immobile and the fractures do not change their positions.
@@ 2.1 Governing equations for the flow
In this section, we briefly describe the PDEs and refer to [36, 38] for details. From the modeling point of view, we consider the porous medium $\mathscr{M}\subset\mathcal{D}$ surrounding the fracture with the porosity $\phi_{m}:\mathscr{M}\to(0,1)$ and the permeability ${\mathbf{K}}_{m}:\mathscr{M}\to\mathbb{R}^{d\times d}$. The fluid phase in this subdomain is characterized by salt mass fraction $c_{m}(t,\mathbf{x}):[0,+\infty)\times\mathscr{M}\to[0,1]$ and pressure $p_{m}(t,\mathbf{x}):[0,+\infty)\times\mathscr{M}\to\mathbb{R}$. So, the flow in $\mathscr{M}$ obeys the standard mass conservation laws for the entire liquid phase and the salt
@@
$\left.\begin{array}[]{l}\partial_{t}(\phi_{m}\rho_{m})+\nabla\cdot(\rho_{m}{\mathbf{q}}_{m})=0\\ \partial_{t}(\phi_{m}\rho_{m}c_{m})+\nabla\cdot(\rho_{m}c_{m}{\mathbf{q}}_{m}-\rho_{m}{\mathbf{D}}_{m}\nabla c_{m})=0\end{array}\right\}\quad x\in\mathscr{M}$ (1)
with the Darcy’s law for the velocity:
${\mathbf{q}}_{m}=-\frac{{\mathbf{K}}_{m}}{\mu_{m}}(\nabla p_{m}-\rho_{m}{\mathbf{g}}),\qquad x\in\mathscr{M},$ (2)
@@
where $\rho_{m}=\rho(c_{m})$ and $\mu_{m}=\mu(c_{m})$ indicate the density and the viscosity of the liquid phase, ${\mathbf{D}}_{m}(t,\mathbf{x}):[0,+\infty)\times\mathscr{M}\to\mathbb{R}^{d\times d}$ denotes the molecular diffusion and mechanical dispersion tensor.
@@
The fracture is assumed to be filled with a porous medium, too. In particular, we assume the validity of the Darcy’s law inside the fracture. Without loss of generality, we consider in this work only one fracture, see [36, 38] for the generalization. The fracture is represented by a surface $\mathscr{S}\subset\mathcal{D}$, $\mathscr{M}\cup\mathscr{S}=\mathcal{D}$, $\mathscr{M}\cap\mathscr{S}=\emptyset$. However, we distinguish between the two sides of the fracture, $\mathscr{S}^{(1)}$ and $\mathscr{S}^{(2)}$, that geometrically coincide with $\mathscr{S}$ but are virtually interfaces of the fracture with $\mathscr{M}$ and have opposite normals $\mathbf{n}^{(1)}$ and $\mathbf{n}^{(2)}$.
@@
For the fluid phase inside the fracture, let $c_{f}:[0,+\infty)\times\mathscr{S}\to[0,1]$ and $p_{f}:[0,+\infty)\times\mathscr{S}\to\mathbb{R}$ be the salt mass fraction and the pressure averaged along the vertical, cf. [21]. Note that in general, the values of $c_{f}$ and $p_{f}$ are not equal to the limits of $c_{m}$ and $p_{m}$ at $\mathscr{S}$. Furthermore, the latter limits may be different on $\mathscr{S}^{(1)}$ and $\mathscr{S}^{(2)}$. We denote them by $c_{m}^{(1)}$ and $c_{m}^{(2)}$, as well as $p_{m}^{(1)}$ and $p_{m}^{(2)}$ as functions on $\mathscr{S}^{(1)}$ and $\mathscr{S}^{(2)}$, respectively. Two sets of the equations are stated for the fracture: The laws that describe the flow along the fracture and the ones modeling the mass exchange through the interfaces $\mathscr{S}^{(1)}$ and $\mathscr{S}^{(2)}$.
@@
For the lateral flow in the fracture, we impose the analogues of the mass conservation laws (1):
@@
where $\nabla^{\mathscr{S}}$ denotes the differential operators on the manifold $\mathscr{S}$ and $\epsilon$ the fracture width. The Darcy velocity along the fracture is
@@
In (34), $\rho_{f}=\rho(c_{f})$ and $\mu_{f}=\mu(c_{f})$ are the density and the viscosity of the fluid phase in the fracture. Furthermore, $\phi_{f}:\mathscr{S}\to[0,1]$ denotes the porosity associated with the material in the fracture, $K_{f}:\mathscr{S}\to\mathbb{R}$ its lateral permeability and $D_{f}(t,x):[0,+\infty)\times\mathscr{S}\to\mathbb{R}$ the lateral diffusion and dispersion tensor in the fracture.
@@
In (6), $K_{fn}^{(k)}:\mathscr{S}^{(k)}\to\mathbb{R}$ is the normal permeability and $D_{fn}^{(k)}:[0,+\infty)\times\mathscr{S}^{(k)}\to\mathbb{R}$ is the normal diffusion and dispersion coefficient of the fracture – bulk medium interface $\mathscr{S}^{(k)}$.
@@
Furthermore, for the inner edges of fractures, we impose the no-flux boundary conditions so that no mass exchange is possible there, cf. [20, 21].
Equations (1), (3) and (7) with the velocities (2), (4) and (6) as well as particular specifications of the parameters $\rho$, $\mu$, $\phi_{m,f}$, ${\mathbf{K}}_{m}$, ${\mathbf{D}}_{m}$, $K_{f}$, $K_{fn}$, $D_{f}$ and $D_{fn}$ form the complete model for the density-driven flow in the porous medium. This system must be closed by boundary conditions for $c_{m}$, $c_{f}$, $p_{m}$ and $p_{f}$ on $\partial\mathcal{D}$ as well as initial conditions for $c_{m}$ and $c_{f}$ at $t=0$.

### v3

@@ Abstract
We use the Multi Level Monte Carlo method to estimate uncertainties in a Henry-like salt water intrusion problem with a fracture. The flow is induced by the variation of the density of the fluid phase, which depends on the mass fraction of salt. While the fracture’s location is fixed, its aperture is uncertain. In our setting, porosity and permeability vary spatially and recharge is time-dependent. So we introduce three random variables, one controlling both the porosity and permeability fields, one for the fracture width and one for the intensity of recharge. For each realization of these uncertain parameters, the evolution of mass fraction and pressure fields is modeled using a system of non-linear, time-dependent PDEs with a solution discontinuity at the fracture. These uncertainties propagate, affecting the distribution of salt concentration, a key factor in water resource quality. We show that the MLMC method can be successfully applied to this problem. It significantly reduces the computational cost compared to classical Monte Carlo methods by effectively balancing discretisation and statistical errors, and by evaluating multiple scenarios over different spatial and temporal mesh levels. The deterministic PDE solver, using the ug4 library, runs in parallel to compute all stochastic scenarios.
@@ 1 Introduction
Density-driven flow is a phenomenon observed in several natural and engineering contexts, including weather patterns, atmospheric and oceanic circulation, geothermal energy extraction, and cooling of electronic and nuclear reactors. In hydrogeology, modelling density-driven subsurface flow and salt transport in fractured aquifers presents significant challenges. These arise from the complex, parameter-sensitive flow patterns caused by fractures and various uncertain parameters. Fractures introduce heterogeneity into the aquifer, affecting fluid flow and making it difficult to predict key properties. In addition, fractures can become sealed over time, reducing their permeability and further affecting fluid flow.
Uncertainties in fracture geometry, porosity, permeability and recharge have a significant impact on the prediction of the evolution of subsurface salinity (cf. [1, 3, 4, 37, 47]). In real reservoirs, the number, location and aperture of fractures are inherently uncertain. One possible solution is to model these uncertainties, build a surrogate model and perform sensitivity analysis (see [6, 48]) to identify critical uncertainties and assess their impact on the solution. However, this approach is complex.
The combination of MLMC with the geometric multigrid method is a promising strategy. The geometric multigrid method (GMG) uses a hierarchy of grids and Multi-Level Monte Carlo (MLMC) involves sampling from different grids. The multigrid method is already implemented in our UG4 software library and in many other codes. We note that both MLMC and surrogate models can be implemented in an almost black-box fashion.
In our previous work [35], we applied the MLMC method to a similar problem, albeit without a fracture in the aquifer. The problem addressed in this paper is more complex, involving a greater number of equations, additional uncertainties, and a discontinuous solution. In other related works [33] and [34], the authors employed the Polynomial Chaos Expansion (PCE) method to solve the Elder problem, another density-driven flow problem. However, the Elder problem differs significantly from the current problem. While the authors successfully applied PCE, they encountered certain limitations, and it remains unclear whether their methods and insights can be effectively extended to the present problem with a fractured reservoir. Multi-level multi-fidelity surrogate models have been explored in [8].
In [43], the authors are solving the seawater intrusion problem under multiple sources of uncertainty. They use the generalised polynomial chaos expansion approximation (gPCE) to investigate how incomplete knowledge of system properties affects the assessment of global quantities. They consider very different input uncertainties, such as the gravity number, the permeability anisotropy ratio, and the transverse and longitudinal Peclet numbers. Their uncertainties are scalar values, whereas our uncertainties include random fields (porosity and permeability) and random process (recharge).
@@
In [21, 22, 42], the authors modeled and computed the density-driven flow in fractured porous media with deterministic settings. The fractures considered were represented by a $(d-1)$-dimensional manifolds in a $d$-dimensional domain. They developed a special numerical technique based on the finite volume method and implemented it in the ug3 library. In our present work, we apply the same model, the same discretization and solvers to compute the multiple uncertain scenarios. Other methods based on $(d-1)$-dimensional representation of fractures were considered in [1, 3, 4, 37, 47, 44, 20]. It is worth to remark that this type of approaches with the low-dimensional representation of the fractures becomes inaccurate for large fracture apertures. As demonstrated in [22] where simulation results of similar deterministic settings with $d$- and $(d-1)$-dimensional representations of the fractures are compared, this issue becomes significant for the density-driven flow: Due to the non-linearity of the equations, vortices arising in thick fractures change the flow and transport in the entire domain essentially. Taking this observation into account, in the present work, we restrict ourselves to sufficiently thin fractures.
@@
Many techniques can be used to quantify uncertainties. One classical method is Monte Carlo (MC) sampling. It has a well-known disadvantage — a slow convergence $\mathcal{O}(\frac{1}{\sqrt{N}})$. Other relative recent techniques such as surrogate models and stochastic collocation may require a few hundred time-consuming simulations and assume a certain smoothness of the quantity of interest (QoI). Another class of methods is the class of perturbation methods [12]. The idea is to decompose the QoI with respect to random parameters in a Taylor series. The higher order terms can be neglected for small perturbations, simplifying the analysis and numerics. These methods assume that the random perturbations are small. For larger perturbations, these methods usually do not work.
Recent efforts have benchmarked various discrete fracture models, as detailed in [15] and [5], albeit in deterministic settings rather than uncertain ones. In [15], several benchmarks were presented for numerical schemes addressing single-phase fluid flow in fractured porous media. These benchmarks compared multiple solution strategies and progressively increased the complexity of network geometry—such as intersecting fractures—and physical parameters, including low and high fracture-matrix permeability ratios and heterogeneous fracture permeabilities. Similarly, [5] introduced four benchmark cases for single-phase flow in three-dimensional fractured porous media. These cases were meticulously designed to test the capability of various methods to manage the complexities inherent in the geometric structures of fracture networks. A total of 17 numerical methods were collected and compared.
As a model problem, we consider the benchmark from [21] which is a generalization of the Henry problem first introduced in [27]. The Henry problem became a benchmark for numerical solvers for the density-driven groundwater flow (see [51, 46, 45, 14].
Novelty. We combined the geometric multi-grid and MLMC methods to quantify uncertainties in a discontinuous, time-dependent, nonlinear, density-driven flow problem characterized by a system of coupled partial differential equations. The primary objective was to apply and to test the MLMC method. Both the MLMC method and parallel geometric multigrid methods were employed to estimate the mean and variance of the salt mass fraction. The computational domain encompasses a fracture. Three parameters —– porosity, recharge and fracture aperture — are treated as uncertain. The porosity depends on the spatial coordinates, the recharge is time-dependent, and the fracture aperture is modeled by a random variable. The permeability of the medium is a function of its porosity. Our numerical experiments confirm that, under certain settings, the MLMC method is applicable and can compute the mean up to 100 times faster than standard Monte Carlo methods.
@@
This work is structured as follows. Section 2 describes the model. Section 3 provides a description of the numerical methods, with a review of the well-known MLMC method in Section 3.2. Section 4 presents the numerical results, including the analysis of the model problem, computation of various statistics, and the performance evaluation of both the MLMC method and the parallel multigrid solver. Finally, we conclude this work with a discussion in Section 5.
@@ 2 Modeling and Problem Settings
We examine the density-driven flow of a liquid phase (salt solution in water) within a fractured aquifer, where the salt mass fraction varies. Fractures in the aquifer are geological formations with a negligible size in one geometric direction compared to the domain’s scales, but the permeability of their filling material is significantly higher than that of the surrounding medium. This disparity presents substantial challenges when treating fractures as full-dimensional subdomains in simulations. Specifically, resolving these features requires very fine grids, and the sharp contrast in model parameters drastically reduces numerical method efficiency. Consequently, several approaches have been developed to address these issues (cf. [1, 4, 17, 13, 36, 30, 42]).
A notable advantage of these models for sampling methods is the ability to alter the fracture width without remeshing the entire domain. In this study, we assume the porous matrix remains immobile, and the fractures retain their positions.
@@ 2.1 Governing equations for the flow
In this section, we briefly describe the partial differential equations (PDEs) of the model. For further details, we refer to [40, 42]. We denote by $\mathcal{D}\subseteq\mathbb{R}^{d}$ the entire domain representing the aquifer with $\mathscr{M}\subseteq\mathscr{D}$ being its part surrounding the fracture. The porosity of this part is $\phi_{m}:\mathscr{M}\to(0,1)$ and its permeability ${\mathbf{K}}_{m}:\mathscr{M}\to\mathbb{R}^{d\times d}$.
The fluid phase in $\mathscr{M}$ is characterized by the salt mass fraction $c_{m}(t,\mathbf{x}):[0,+\infty)\times\mathscr{M}\to[0,1]$ (where $1$ corresponds to the saturated brine) and pressure $p_{m}(t,\mathbf{x}):[0,+\infty)\times\mathscr{M}\to\mathbb{R}$. Thus, the flow and transport in $\mathscr{M}$ obey the standard conservation of mass laws for the entire liquid phase and the salt
@@
$\left.\begin{array}[]{l}\partial_{t}(\phi_{m}\rho_{m})+\nabla\cdot(\rho_{m}{\mathbf{q}}_{m})=0\\ \partial_{t}(\phi_{m}\rho_{m}c_{m})+\nabla\cdot(\rho_{m}c_{m}{\mathbf{q}}_{m}-\rho_{m}{\mathbf{D}}_{m}\nabla c_{m})=0\end{array}\right\}\quad x\in\mathscr{M}.$ (1)
In (1), we assume the Darcy’s law for the velocity:
${\mathbf{q}}_{m}=-\frac{{\mathbf{K}}_{m}}{\mu_{m}}(\nabla p_{m}-\rho_{m}{\mathbf{g}}),\qquad x\in\mathscr{M}.$ (2)
@@
In (12), $\rho_{m}=\rho(c_{m})$ and $\mu_{m}$ are the density and the viscosity of the liquid phase, ${\mathbf{D}}_{m}(t,\mathbf{x},{\mathbf{q}}_{m})$ denotes the diffusion and mechanical dispersion tensor.
@@
The fractures are also considered to be filled with a porous medium. In particular, we assume Darcy’s law inside the fractures. Without loss of generality, we consider only one fracture in this work, see [40, 42] for a generalization. The fracture is represented by a surface $\mathscr{S}\subset\mathcal{D}$, ¯$\overline{\mathscr{M}}\cup\overline{\mathscr{S}}=\overline{\mathcal{D}}$, $\mathscr{M}\cap\mathscr{S}=\emptyset$. However, we distinguish between the two sides of the fracture, $\mathscr{S}^{(1)}$ and $\mathscr{S}^{(2)}$, which geometrically coincide with $\mathscr{S}$, but virtually are interfaces of the fracture with $\mathscr{M}$ and have opposite normals $\mathbf{n}^{(1)}$ and $\mathbf{n}^{(2)}$.
@@
For the fluid phase inside the fracture, let $c_{f}:[0,+\infty)\times\mathscr{S}\to[0,1]$ and $p_{f}:[0,+\infty]\times\mathscr{S}\to\mathbb{R}$ are the salt mass fraction and the pressure averaged along the vertical, cf. [22]. Note that in general the values of $c_{f}$ and $p_{f}$ are not equal to the limits of $c_{m}$ and $p_{m}$ at $\mathscr{S}$. Furthermore, the latter limits may be different on $\mathscr{S}^{(1)}$ and $\mathscr{S}^{(2)}$. We denote them by $c_{m}^{(1)}$ and $c_{m}^{(2)}$, as well as $p_{m}^{(1)}$ and $p_{m}^{(2)}$ as functions on $\mathscr{S}^{(1)}$ and $\mathscr{S}^{(2)}$, respectively. Two sets of equations are given for the fracture: the laws describing the flow along the fracture and those modeling the mass exchange through the interfaces $\mathscr{S}^{(1)}$ and $\mathscr{S}^{(2)}$.
@@
For the flow and transport along the fracture, we impose the analogues of the mass conservation laws (1):
@@
where $\nabla^{\mathscr{S}}$ denotes the differential operators on the manifold $\mathscr{S}$ and $\epsilon$ the fracture width. The tangential Darcy’s velocity in the fracture is
@@
In (34), $\rho_{f}=\rho(c_{f})$ and $\mu_{f}=\mu(c_{f})$ are the density and the viscosity of the fluid phase in the fracture. Furthermore, $\phi_{f}:\mathscr{S}\to(0,1)$ denotes the porosity associated with the material in the fracture, $K_{f}:\mathscr{S}\to\mathbb{R}$ its tangential permeability and $D_{f}(t,x):[0,+\infty)\times\mathscr{S}\to\mathbb{R}$ the tangential effective diffusion and dispersion tensor in the fracture.
@@
In (6), $K_{fn}^{(k)}:\mathscr{S}^{(k)}\to\mathbb{R}$ is the normal permeability and $D_{fn}^{(k)}:[0,+\infty)\times\mathscr{S}^{(k)}\to\mathbb{R}$ is the normal effective diffusion and dispersion coefficient of the fracture – bulk medium interface $\mathscr{S}^{(k)}$.
@@
Furthermore, for the inner edges (in 3d) or the inner ends (in 2d) of the fractures, we impose the no-flux boundary conditions so that no mass exchange is possible there, cf. [21, 22].
Equations (1), (3) and (7) with the velocities (2), (4) and (6) as well as particular specifications of the parameters $\rho$, $\mu$, $\phi_{m,f}$, ${\mathbf{K}}_{m}$, ${\mathbf{D}}_{m}$, $K_{f}$, $K_{fn}$, $D_{f}$ and $D_{fn}$ form the complete model for the density-driven flow in the porous medium with the fracture. This system must be closed by boundary conditions for $c_{m}$, $c_{f}$, $p_{m}$ and $p_{f}$ on $\partial\mathcal{D}$ as well as initial conditions for $c_{m}$ and $c_{f}$ at $t=0$.